OpenStudy (anonymous):

how do you find if: f(x)=2x-1, g(x)=x^2+1, h(x)=5-3x...find (f+g)(x)

OpenStudy (unklerhaukus):

\[(f+g)(x)=f(x)+g(x)=2x-1+x^2+1=\cdots\]

OpenStudy (anonymous):

how do i combine them again?

OpenStudy (unklerhaukus):

well the two constant terms can be simplified

OpenStudy (anonymous):

yep i got that, the answer i got was \[2x+x^{2}\] is this right?

OpenStudy (unklerhaukus):

yep that is Right \(\checkmark\)

OpenStudy (anonymous):

yay! okay, now what if its subtracting (f-g)(x)

OpenStudy (unklerhaukus):

\[(f-g)(x)=f(x)-g(x)=\cdots\]

OpenStudy (unklerhaukus):

\[f\circ g(x)=f(g(x))\]

OpenStudy (anonymous):

hahah i get that, I mean i get confused with all the minus- so (2x-1)-(x^2+1)

OpenStudy (unklerhaukus):

your right so far, now expand the brackets and remember to distribute the negative for the term on the right

OpenStudy (anonymous):

2x-2-x^2.. thats what i got and I dont think thats right

OpenStudy (unklerhaukus):

that is right \[(f-g)(x)=2x-2-x^2\qquad\checkmark\]

OpenStudy (anonymous):

yes!! im actually understanding, okay Im having a little trouble with (f/g)(x)

OpenStudy (unklerhaukus):

\[ f(x)=2x-1;\qquad g(x)=x^2+1\] \[(f/g)(x)=f(x)\div g(x)=\frac{2x-1}{x^2+1}=\cdots\]

OpenStudy (anonymous):

okay i got that far, what do i do after i cancel out the -1 and the +1

OpenStudy (unklerhaukus):

the only way to simplify this is to break the fraction if to two components like this \[\frac{a+b}{c}=\frac ac+\frac bc\]

OpenStudy (unklerhaukus):

you can only cancel Factors common to the numerator and denominator in a fraction

OpenStudy (anonymous):

okay so im really confused. I was going to guess 2x/x^2

OpenStudy (unklerhaukus):

\[(f/g)(x)=\frac{2x-1}{x^2+1}=\cdots\] \[\frac{a+b}{c}=\frac ac+\frac bc\] \[a=2x~;\qquad b=-1~;\qquad c=x^2+1\]

OpenStudy (anonymous):

alright! I got that part written down, i just dont know what to do after that

OpenStudy (unklerhaukus):

make the substitution

OpenStudy (anonymous):

can i cancel both +1's?

OpenStudy (unklerhaukus):

no, the 1's stay

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

can i do something with both denominators?

OpenStudy (unklerhaukus):

say if i had \[\frac {2y-3}{y^4+7}\] i could use the substitution\[\frac{a+b}{c}=\frac ac+\frac bc\] to arrive at\[\frac {2y-3}{y^4+7}=\frac {2y}{y^4+7}-\frac3{y^4+7}\] this cannot be simplified further

OpenStudy (anonymous):

1x/2x^2+2... or am i completely wrong?

OpenStudy (unklerhaukus):

your answer should be two fractions

OpenStudy (anonymous):

ohhh so its 2x/x^2+1 - -1/x^2+1

OpenStudy (unklerhaukus):

yes, but with only one negative sign

OpenStudy (anonymous):

yeah thats just my subtract sign in the middle

OpenStudy (unklerhaukus):

if you mean this\[\frac{2x}{x^2+1} - \frac{1}{x^2+1}\]you are correct

OpenStudy (anonymous):

oh I see what your saying!

OpenStudy (anonymous):

will you check one for me to see if i did it right?

OpenStudy (unklerhaukus):

sure

OpenStudy (anonymous):

(f of h) (2x-1)(5-3x)

OpenStudy (anonymous):

is my answer -6x-5?

OpenStudy (unklerhaukus):

thats not right

OpenStudy (anonymous):

-6x^2?

OpenStudy (anonymous):

6x^2+7x-5?

OpenStudy (unklerhaukus):

\[f(x)=2x-1\qquad \qquad h(x)=5-3x\] \[(f \circ h)=f(h(x))\]\[\qquad =f(5-3x)\]\[\qquad=2(5-3x)-1\]

OpenStudy (anonymous):

When given the question: find f(g(x)), we plug in f(x) into g(x).true or false?

OpenStudy (anonymous):

When given the question: find f(g(x)), we plug in f(x) into g(x).true or false?

OpenStudy (anonymous):

shoot i got this mixed up with multiplication...

OpenStudy (unklerhaukus):

not quite, we find g(x) and then find f(g(x))

OpenStudy (anonymous):

stop NOELL! im getting help

OpenStudy (anonymous):

shes my sister...

OpenStudy (unklerhaukus):

the product of functions is\[f\times g(x)=f(x)\times g(x)\] _________ \[f\circ g(x)=f(g(x))\] is called the composition of functions

OpenStudy (unklerhaukus):

?

OpenStudy (unklerhaukus):

hello

OpenStudy (anonymous):

noniebug is my sister. sorry! but okay back to me! would the answer be 6x-8?

OpenStudy (unklerhaukus):

that is \(h\circ f\)

OpenStudy (unklerhaukus):

\[h\circ f\neq f\circ h\]

OpenStudy (anonymous):

8-6x

OpenStudy (unklerhaukus):

nope

OpenStudy (anonymous):

uhhh im confused then.

OpenStudy (unklerhaukus):

\[(f \circ h)=f(h(x)) =f(5-3x) =2(5-3x)-1=\cdots\]

OpenStudy (anonymous):

do i put the 2 with everything?

OpenStudy (unklerhaukus):

\[2(5−3x)−1=(2\times 5)+(2\times-3x)-1\]

OpenStudy (anonymous):

10-6x-1? I dont know

OpenStudy (unklerhaukus):

thats good, now simplify the constants

OpenStudy (anonymous):

9-6x

OpenStudy (unklerhaukus):

thats is correct \(\checkmark\)

OpenStudy (anonymous):

yay! okay so i see where i got off at, i distubuted the 2 with the -2

OpenStudy (anonymous):

-1**

OpenStudy (anonymous):

now i have a multiplication one hah, i skipped it on acciendent. (f x g)(x)

OpenStudy (unklerhaukus):

you distributed the 2 across onto the 5 and the -3x

OpenStudy (unklerhaukus):

\[f\times g(x)=f(x)\times g(x)\]

OpenStudy (anonymous):

yeah but i went wrong because i kept distributing all the way to the -1 and i wasn't suppose to!

OpenStudy (unklerhaukus):

ah i see

OpenStudy (anonymous):

but okay! (2x-1)(x^2+1)

OpenStudy (unklerhaukus):

yep

OpenStudy (anonymous):

i get confused when mulitpling is it 2x^3 or 3x^2

OpenStudy (anonymous):

just the first step?

OpenStudy (unklerhaukus):

\[2x\times x^2=2x^3\]

OpenStudy (anonymous):

so is the answer 2x^3+2x-x^2

OpenStudy (unklerhaukus):

there is one more term

OpenStudy (anonymous):

2x^3+2x-1x^2?

OpenStudy (unklerhaukus):

\[(2x-1)(x^2+1)=2x\times x^2+2x-x^2-1\]

OpenStudy (anonymous):

2x+2x-1? the x^2"s cancel out?

OpenStudy (anonymous):

which would make it 4x^2-1 or am I wrong?

OpenStudy (unklerhaukus):

no remember \(2x×x^2=2x^3\)

OpenStudy (anonymous):

2x^3+2x-x^2-1..I dont know after that :(

OpenStudy (unklerhaukus):

thats is as simplified as you can get.

OpenStudy (anonymous):

oh so thats it??? oh i tried to cancel the 1's out when its multiplication...ha

OpenStudy (unklerhaukus):

that is it

OpenStudy (anonymous):

yipee! okay do you have time for one more? its like the (f of g) one..

OpenStudy (unklerhaukus):

if your quick,

OpenStudy (anonymous):

kay (g of h)

OpenStudy (anonymous):

g=x^2+1 and h=5-3x

OpenStudy (unklerhaukus):

\[g\circ h(x)=g(h(x))=\] substitute h in

OpenStudy (anonymous):

is it 5x^2-3x^3+1?

OpenStudy (unklerhaukus):

not quite \[g(5-3x)=(5-3x)^2+1 \]

OpenStudy (anonymous):

..3x^2+26???????

OpenStudy (unklerhaukus):

not right \[(a+b)^2=(a+b)(a+b)\]\[=a\times a+a\times b+b\times a+b\times b\]\[=a^2+2ab+b^2\]

OpenStudy (anonymous):

9x^2+25?

OpenStudy (anonymous):

9x^2+26